OpenStudy (anonymous):

When Amy was shopping at Gabes's Grocery, she saw a sign for soup that advertised 4 cans of soup for $5. She looked in the sales flyer of another grocery store called BIG DEALS. She saw that this week, she could buy one can of the same soup for $3 and get another can for free. Which is the better deal?

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

you want "price per can" which means \[\frac{\text{dollars}}{\text{can}}\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[\frac{5}{4}=1.25\] \[\frac{3}{2}=1.5\] so first one is cheaper

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So big deals?